For the line going through (3,-1) and (-2,4) write the equation in point-slope form, convert to slope-intercept form and find the x-intercept and y-intercept.
Answer by ankor@dixie-net.com(22740)

For the line going through (3,-1) and (-2,4)
write the equation in point-slope form,
y - y1 = m(x - x1)
-1-4 = m(3 -(-2))
-5 = m(3+2)
-5 = 5m
m = -1
y - 4 = -1(x + 2)
convert to slope-intercept form and
y - 4 = -x - 2
y = -x - 2 + 4
y = -x + 2
find the x-intercept and y-intercept.
x intercept when y=0
-x + 2 = 0
2 = x
y intercept when x=0
y = 0 + 2
y = 2